OpenStudy (rmrjr22):

Need Urgent Help!! Let the Parametric curve; x = 8sin(t) y = 4Sqroot(3)Cos(t) 0

OpenStudy (tkhunny):

Have you considered finding the value of t that puts you at (4,6)? Don't forget to make sure that you get there only once.

OpenStudy (rmrjr22):

making a chart and plugging every point in until i reach (4,6)?

OpenStudy (tkhunny):

You're not serious! Why not solve for t in \(8\sin(t) = 4\), finding all solutions in \(0 < t < 3\pi\).

OpenStudy (rmrjr22):

thats what im trying to figure out. Going blank on parametric equations

OpenStudy (rmrjr22):

so ur saying the arcsin(1/2) = t?

OpenStudy (rmrjr22):

then plug in t into the y equation?

OpenStudy (tkhunny):

That is not a parametric equation. It is a simple trig equation. You should be up to it. \(\sin(t) = \dfrac{1}{2}\) \(t = \dfrac{\pi}{6},\dfrac{5\pi}{6},\dfrac{13\pi}{6},\dfrac{17\pi}{6}\) Please stop "plugging in". We hope you will learn to call it "substitution".

OpenStudy (tkhunny):

No, do NOT use the arcsine function. It has a range of only \(-\dfrac{\pi}{2} < t < \dfrac{\pi}{2}\). You need \(0 < t < 3\pi\). It's just trigonometry.

OpenStudy (rmrjr22):

arent we trying to find t so we can use it and in this case would we have to find the arcsin of 1/2 to get (pi)/6?

OpenStudy (tkhunny):

If you use the arcsine function, your ONLY possible solution will be \(\pi/6\). There are three more possible solutions. I listed them above.

OpenStudy (tkhunny):

What we just demonstrated is that x takes on the value 4 a total of four times in \(0 < t < 3\pi\). Now, we have to hope that y takes on the value 6 for at least one of those values of t.

OpenStudy (rmrjr22):

y = 6 at (pi)/6

OpenStudy (tkhunny):

\( 6 = 4\sqrt{3}\cos(t)\) \(\cos(t) = \dfrac{\sqrt{3}}{2}\) \(t = \dfrac{\pi}{6}, \dfrac{11\pi}{6}, \dfrac{13\pi}{6}\) Whoops! if you just go with \(\pi/6\), you WILL MISS the other solution.

OpenStudy (tkhunny):

You must be thorough. Don't just guess or hope that you have it because you managed to trip over something interesting.

OpenStudy (rmrjr22):

ok, so if they give me a point. Respectfully substitute x or y and solve for t. Correct?

OpenStudy (tkhunny):

I'm a little troubled by the original problem statement. There are TWO solutions to this problem and it asks for the slope of "the tangent line". There are TWO such tangent lines. You must be sure that you find ALL solutions. I suspect, because the problem statement goes all the way to \(3\pi\), that it was intended that you would find two solutions. The problem statement is worded oddly. You'll be one of few who finds both solutions.

OpenStudy (rmrjr22):

ok, so it wasnt just me. I was really confused. thought there was only supposed to be one solution. Thank you